OpenStudy (anonymous):

Evaluate the following expression using the values given. Find 3x - y - 3z if x = -2, y = 1, and z = -2. A.) -13 B.) -1 C.)1 D.)13

OpenStudy (kropot72):

Sustitute the given values in the expressio as follows: (3*-2) - (1) - (3*-2) = -6 - 1 +6 = -1 So the correct choice is B